askew616:

What is the equation of the line that passes through the points (−2, 3) and (2, 7)? (5 points)

askew616:

There is no 0 answer choice

ILikeToEatChildren:

@aiannafloyd wrote:
The equation of the line that passes through the points (-2, 3) and (2, 7) is x - y + 5 = 0.
no direct answers, you should explain it to her

askew616:

The choices are x − y = −1 x − y = −2 x − y = − 5 x − y = − 6

askew616:

Can you plz help me? The answer is not 0 it is not a choice

askew616:

Does anyone agree with -2

askew616:

This is very important anyone have answers

askew616:

Okay thanks guys

AiannaFLOYD:

You´re welcome!

askew616:

Are you guys always on here

RAiNB0WGUTZZ:

Yeah we are

AiannaFLOYD:

Yes, I am all day

RAiNB0WGUTZZ:

Well I am

askew616:

How

askew616:

where are you guys from

RAiNB0WGUTZZ:

Japan

AiannaFLOYD:

Im from Cali, but live in Oklahoma

askew616:

I live in Tennessee

askew616:

How do you fan someone

RAiNB0WGUTZZ:

Press their username and it'll pop up

RAiNB0WGUTZZ:

Or just hover over it

carmelle:

I politely disagree with the answer you have chosen, B is not the correct answer. A simple way to get the answer is to plug in the values and see what they subtract to... (x,y) -> (-2,3) (x,y) -> (2,7) -2 - 3 = ? 2 - 7 = ? What's the answer to these two equations?